Publishing With Maple

Douglas B. Meade

Department of Mathematics University of South Carolina

Columbia, SC 29208

      meade@math.sc.edu
http://www.math.sc.edu/˜ meade/

Philip Yasskin

Department of Mathematics Texas A & M University College Station, TX 77843

    yasskin@math.tamu.edu
http://www.math.tamu.edu/˜ philip.yasskin/

18 November 2000

November 2000  ICTCM XIII (Atlanta)  0

Publishing With Maple

Douglas B. Meade

Department of Mathematics University of South Carolina

Columbia, SC 29208

      meade@math.sc.edu
http://www.math.sc.edu/˜ meade/

Philip Yasskin

Department of Mathematics Texas A & M University College Station, TX 77843

    yasskin@math.tamu.edu
http://www.math.tamu.edu/˜ philip.yasskin/

18 November 2000

November 2000  ICTCM XIII (Atlanta)  1

Publishing with Maple

\underline{Panelists}

Eugene Herman Grinnell College
Tom Lee  Waterloo Maple
Robert Lopez  Rose Hulman

Institute of Technology

Marlene Thom Addison Wesley Longman
Joe Vetere  Addison Wesley Longman

\underline{Organizers}

Douglas Meade University of South Carolina
Philip Yasskin Texas A & M University

Basic Questions

1. Describe, briefly, your most significant publishing experience with Maple.

2. What aspects of the publishing process were made simpler because of Maple?

3. What aspects of the publishing process were made more difficult because of Maple?

4. What would you do differently if you were starting a project today?

5. What changes/additions to Maple would improve the publishing process?

6. What changes in the publisher’s procedures would improve the publishing process (for you)?

November 2000  ICTCM XIII (Atlanta)  3

Closing Comments

Suppose I am considering a new publishing project with Maple.

1. What advice would you give?

2. Where could they get support for the project?

3. Where should they look for additional information?